OpenStudy (anonymous):

write equation in vertex form? y=x^2-10x+5

OpenStudy (anonymous):

x^2-10x___+5 = 0 move 5 to the side x^2-10x=25____+5-25 = 0 take half of b and square it, add to the equation subtract from the end. (x-5)^2 - 20 = 0 where vertex is point (5,-20)